OpenStudy (anonymous):

PLZ HELP WILL MEDAL !!!!!!!!!!!!!!!!!!!!!!!! Find the number of hours of rental for which the cost would be the same to rent from either surf shop. Surf and Sand Surf Shop rents surfboards for a $25 flat rate, plus $2.50 per hour. Bikini Bottom Surf Shop rents surfboards for a $10 flat rate, plus $4.00 per hour.

OpenStudy (anonymous):

@bibby

OpenStudy (anonymous):

at 10 hours the rental costs are the same

OpenStudy (anonymous):

how????

OpenStudy (anonymous):

y=2.5(x)+25 x=number of hours, 25 represents flat rate cost y=4(x)+10 x= number of hours, 10 represents flat rate cost

OpenStudy (anonymous):

oh ok thx